OpenStudy (firejay5):

Identify the property that justifies the statement: If 4x - 3 = 7, then 4x = 10. -----> Is this substitution property?

hero (hero):

Are you familiar with the properties of equality such as Addition Property of Equality or Subtraction Property of Equality?

OpenStudy (firejay5):

yes

hero (hero):

So which of those apply here? When you perform the following operation: \[4x - 3 + 3 = 7 + 3\] Which is adding 3 to both sides, what property is that?

OpenStudy (firejay5):

4x -3 = addition property

hero (hero):

@Firejay5, 4x - 3 is not equal to the addition property. The operation that you perform to get from 4x - 3 = 7 to 4x = 10 uses the addition property of equality.

hero (hero):

So which of those apply here? When you perform the following operation: \[4x - 3 + 3 = 7 + 3\] Which is adding 3 to both sides, what property is that?

hero (hero):

Are you familiar with the properties of equality such as Addition Property of Equality or Subtraction Property of Equality?

OpenStudy (firejay5):

yes

OpenStudy (firejay5):

Isn't it substitution property

OpenStudy (firejay5):

@Hero Is it not substitution property

hero (hero):

No, it is not substitution property. You're confusing yourself. I know why you think it is substitution property. We haven't solved for x yet.

hero (hero):

We're only trying to figure out the property that is used to get from 4x - 3 = 7 to 4x = 10

hero (hero):

Trust me, there is no substitution property involved with adding 3 to both sides.

hero (hero):

We're only trying to figure out the property that is used to get from 4x - 3 = 7 to 4x = 10

OpenStudy (firejay5):

It's saying 4x = 10, replace 4x to 10 to get 10 - 3 = 7

hero (hero):

Negative. That's not what it's saying. I explained it to you already. At this point, you're only confusing yourself.

hero (hero):

Are you familiar with the properties of equality such as Addition Property of Equality or Subtraction Property of Equality?

OpenStudy (firejay5):

10 - 3 = 7 is that not true

hero (hero):

It's true, but substitution property has nothing to do with the question.

OpenStudy (firejay5):

so you are trying to add 3 to both sides to get 10 and then divide 4x by 10

hero (hero):

@Firejay5, I'm only going to show you this once more: